OpenStudy (anonymous):

A small class has 9 students, 4 of whom are girls and 5 of whom are boys. The teacher is going to choose two of the students at random. What is the probability that the first student chosen will be a girll be a boy? Write and the second wilyour answer as a fraction in simplest form.

OpenStudy (anonymous):

girl = 4/9 boy= 5/9

OpenStudy (anonymous):

most likely a boy

OpenStudy (anonymous):

__ (just choose the one thats most)